All ISS systems continue to function nominally, except as noted previously or below. Day 131 in space for Expedition Five (of a planned 167-day stay).
The third EVA was successfully accomplished in 6hrs 36 min.
After wake-up at 3:46am EDT, the crew found thanks from the ground for assisting in breaking loose of the lower launch restraints on the Airlock (AL) rack, which had prevented the installation of the rack pivot pin.
The spacewalk began at 10:11am EDT, 30 min ahead of schedule, when Dave Wolf and Piers Sellers switched their EMU suits to battery power. Its objectives were (1) connecting ammonia (NH3) jumpers between S0 and S1; (2) removal of two large keel pin assemblies that held S1 in the Atlantis cargo bay; (3) installation of six final SPDs (spool positioning devices) on ammonia QDs (quick disconnects); and (4) repair or replacement, if necessary, of the backup cable cutter on the SSRMS MBS (mobile base system), which had an incompletely retracted safety bolt from its earlier installation. However, at 10:56am, Wolf and Sellers successfully retracted the troublesome safing bolt, which made replacement of the entire IUA (interface umbilical assembly) unnecessary. The new IUA, now not needed, will be stowed back in the Airlock (AL). All objectives were successfully accomplished. The spacewalk ended at 4:47pm EDT, bringing the total time of all 46 ISS EVAs to 285h 25m.
Prior to EVA-2 on 10/12, Dave Wolf had experienced decreased O2 (oxygen) flow rate during the prebreathe period. Thus, for EVA-3, he was advised of the possibility that there was either a flow restrictionor a leak inthe prebreathe mask or the special T-fitting assembly.For today's EVA prebreathe, in addition to putting on a different mask, he also used a differentspecial T-fitting.At 4:11am, the remote-controlled deployment of the loop A TCS (thermal control system) radiator on the S1 truss was successfully completed. [The concertina-like unfolding of the 75-ft. long panel structure was postponed from last night, but the rotation of the large framework holding the folded radiator panels through a full 90 deg was ground-commanded, such that it is now perpendicular to the S1 truss at which it is attached. The S1 TCS will not be activated until next year. The deployment of one of the three sets of radiators was to test its mechanical operation.]
During checkout and initial commanding of the S1 TRRJ (thermal radiator rotary joint) torque motor resolver, the torque motor was automatically shut down. The failure was repeated twice on string 1, and again on string 2 with the same signature. MCC-H decided to inhibit the FDIR (failure detection, isolation, and recovery) software of the TMR (torque motor resolver) and continue the checkout. The TRRJ was successfully rotated to its first hard stop at 113.6 deg on string 1 and to its second hard stop at 114.9 deg on String 2, before being returned to the 0 deg position.
The TVIS treadmill received a new chassis and other components yesterday, followed by activation and checkout. Ground engineers are awaiting the downlink of the checkout data to evaluate the results of the IFM (in-flight maintenance). [During the R&R, the crew noted that of the four gyro cables comprising the VIS (vibration isolation system), the starboard forward cable was completely separated and the port aft has about five or six (of 19 total) threads of wire cable broken on various strands. The starboard aft cable has maybe one or two wires on various strands broken. Based on agreements reached during an emergency FIT (failure investigation team) meeting, a procedure was sent to the crew to attach a piece of Teflon to the bumpers to keep the TVIS motion from causing metal contact with metal. Also, during checkout of the new TVIS chassis, the gap between the rear deck assembly and belt was found to be wider than the previous configuration. This can pose a potential hazard with loose shoelaces, so a step to double-knot shoelaces and a warning note have been added to the crew's TVIS cue card and exercise reference card with SLD (subject load devices) contingency configuration.]
Atlantis performed the second scheduled reboost of the station this morning. ISS attitude control was handed over to the Orbiter at 6:30am, and the maneuver began at 7:20am and lasted 35 min. Attitude gain was 2.3 s.mi. (3.7 km). Attitude control authority returned to the ISS CMGs with thruster assist (TA) at 8:30am.
To fulfill Russia's contractual obligations to Japan/NASDA, the cosmonauts performed video imaging of the Earth surface with the HDTV (high-definition television) equipment, mounted on a handrail at SM window #7.
Daily routine servicing tasks were performed by FE-2 Sergei Treschev (SOSh life support systems maintenance; IMS inventory update file preparation for downlink), FE-1/SO Peggy Whitson (Lab payloads status checkup), and CDR Valery Korzun (BIO-5 Rasteniya-2 ["Plants-2"] experiment checkup).
Preliminary estimates are that the Shuttle delivered between 30 and 40 lbs of oxygen (O2) to the ISS, as well as 55 lbs of nitrogen (N2).
Yesterday the ISS crew delivered a televised greeting to the Euro-Mediterranean Conference on Child Rights in Morocco under the auspices of UNESCO. Mr. Chakib Laroussi, executive director of Morocco Television, and French cosmonaut Patrick Baudry had requested the crew to address the gathering. Greetings were sent to all children participating in the conference, and all attendees, particularly to Her Royal Highness Princess Lalla Meryem and to the National Observatory of Childhood Rights.
Yesterday at about 8:17am EDT, a current spike in BGA 2B (beta gimbal assembly 2B) was reported. Motor stall was observed at about 8:19am. A backdrive of 20 deg was completed, and BGA 2B is currently back in autotrack.
The crew completed their daily physical exercise program.
Sleep period began tonight at 8:46pm EDT, one hour later than yesterday.